If z=cosθ+i sinθ,prove that

zn + z−n = 2cos(nθ), z^n − z^−n = 2i sin(nθ).


1
Expert's answer
2021-12-26T17:44:07-0500

Use De Moivre's formula


"(\cos\theta+i\sin \theta)^n=\cos(n\theta)+i\sin(n\theta) , n\in \Z"

Then


"z^n=(\cos\theta+i\sin \theta)^n=\cos(n\theta)+i\sin(n\theta)"

"z^{-n}=(\cos\theta+i\sin \theta)^{-n}=\cos(-n\theta)+i\sin(-n\theta)"

"=\cos(n\theta)-i\sin(n\theta)"


"z^n+z^{-n}=\cos(n\theta)+i\sin(n\theta)"

"+\cos(n\theta)-i\sin(n\theta)"

"=2\cos(n\theta), n\in \Z"


"z^n-z^{-n}=\cos(n\theta)+i\sin(n\theta)-"

"-(\cos(n\theta)-i\sin(n\theta))"

"=2i\sin(n\theta), n\in \Z"